Determine the standard form of the equation of the line that contains (8,9) and (-9,-8)
Answer by mananth(16946)

x1 y1 x2 y2
8 9 -9 -8

slope m = (y2-y1)/(x2-x1)
( -8 - 9 )/( -9 - 8 )
( -17 / -17 )
m= 1

Plug value of the slope and point ( 8 , 9 ) in
Y = m x + b
9.00 = 8 + b
b= 9 - -8
b= 17
So the equation will be
Y = 1 x + 17
x-y=-17
m.ananth@hotmail.ca